Calculate the pH of a solutions obtained if 40cm³ of 0.2M of HCL was added to 30cm³ of 0.1M NaOH
1
Expert's answer
2020-06-25T07:51:16-0400

40 cm³ 0.2 M HCl = 8 mmol HCl

30 cm³ 0.1 M NaOH = 3 mmol NaOH

After neutralisation 5 mmol HCl will be remaining in 70 cm³ of solution.

Molarity of HCl in the final solution = 5/70 = 0.071

pH = -log[H+]

= -log(0.071)

= 1.14 Answer